OpenStudy (anonymous):

let f be defined by f(x) = x-4 and g be defined by g(x) = {x^2 - 16 /x+4 ,x ≠ -4 and k , x = -4 } . find k such that f(x) = g(x) for all x.

OpenStudy (anonymous):

x=0 is the only answer I think!

OpenStudy (anonymous):

how ? .. can you solve ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k=0....for x ≠ -4 we already have g(x) = {x^2 - 16 /x+4 =(x-4)*(x+4)/x+4 for x ≠ -4 g(x)=x-4 for x=-4 therefore g(-4)=k=f(-4)=0

OpenStudy (anonymous):

k\[x-4=x ^{2}-16/(x+4) => x-4 = (x ^{3}+4x-16)/x+4 \] \[x ^{2}-16 = x ^{3}+4x-16 => -x ^{3}-4x+x ^{2}=0\] \[-x(x ^{2}+4-x)=0 =>x=0\]

OpenStudy (anonymous):

@mas_gh90 Sorry but, you made a mistake !

OpenStudy (anonymous):

thanks!can u explain more?

OpenStudy (anonymous):

first , the question is to find k not x.

OpenStudy (anonymous):

@mas_gh90 k is the image of -4 by g ! I hope it's clear now !...

OpenStudy (anonymous):

yeah it is!thanks :)

OpenStudy (anonymous):

np :) ;)